# GRU: Gated Recurrent Unit with Spatial Attention for Spatio-Temporal Disaggregation

### v3

@@ 6.2. GRU Improves Neural Results
From Table 3, we observe that our proposed GRU has the best performance among five out of six tasks, delivering an 2.0% improvement over the best neural models on average, and 19.3% improvement over the best classical methods (HR) on average. The largest improvement is observed when we disaggregate from PUMA to BLOCK with NYC Taxi data ā on average, GRU reduces taxi count error per hour in the entire Manhattan region by 134 relative to the GRU model. For statistical significance tests, we use bootstrap sampling for 10,000 iterations (Jurafsky and Martin, 2008; Sonderegger and Buscaglia, 2020): On each iteration, we sample the predictions, then calculate error and compute the proportion that are better than the baseline to provide an empirical estimate of the p-value. We use 0.05 as the significance level. With both the taxi and bikeshare dataset, GRU consistently outperforms GRU with statistically significant improvement (other than NYC Taxi, PUMA to TRACT). Comparing STA with GRU and GRU, we notice that STA models perform poorly on the taxi dataset, even worse than FNN models. With bikeshare data, STA models outperform GRU when disaggregating from PUMA to TRACT and BLOCK, and are less competitive than GRU. Temporal attention is not effective and is incapable of producing consistent results compared with recurrent modeling unit in the disaggregation task. Although attention-based solutions are increasingly popular due to their success in NLP, we find that relatively short sequences and smaller capacity models can still motivate recurrent models.
@@
Figure 4 shows the convergence plots of neural models on two datasets. We notice that other than the disaggregation from PUMA to BLOCK on NYC Bikeshare dataset, GRU models always start with, and maintain, lower error throughout training, thus converging faster than other models. One explanation is that the spatial attention layers in GRU models are aware of the spatial structures due to the containment maps, thus producing better disaggregation results more efficiently.
Ablation Study ā Effectiveness of Containment Maps: We conduct an ablation study to demonstrate the effectiveness of the containment maps. We run GRU models on both datasets without the containment maps in the spatial attention layers. From Table 3, we observe that without the containment maps, the disaggregation results are usually worse than results with the maps. Although the improvement is typically modest, the convergence plots in Figure 4 show that GRU models without maps converge slower than those with the maps, as evidenced by consistently higher error throughout the training. Our interpretation is that the containment maps provide āhintsā during training, steering the learning toward the correct solutions (which always exhibit the containment property), thereby converging faster. Additionally, even without the containment maps, GRU results are still better than GRU on five out of six tasks, indicating the effectiveness of the spatial attention layers, and suggesting that even if the containment property does not hold, these methods will be generally effective.
@@ 6.3. Transfer Learning is Remarkably Effective
We trained GRU models on NYC taxi data and evaluated on bikeshare in the same NYC domain. We fine-tuned with a half day (12 hour-long samples), one day (24), 4 days (96), one week (168 samples), and two weeks (336). Figure 5 shows the results: adding more fine-tuning samples enhances disaggregation performance. Surprsingly, fine-tuning with only one week of data achieves results comparable with, sometimes even better than, the specialized model trained from scratch with bikeshare data. Additionally, fine-tuning a pre-trained model significantly reduces the computational costs. This result has significant implications for practical settings: Training data is not uniformly available for all variables due to technical, privacy, and political effects. But because cities are complex systems, all variables are interdependentĀ (Howe etĀ al., 2022). This result shows that pre-training can capture city dynamics resulting from the built environment and human activity, and that those signals tend to be predictive for multiple variables. We envision developing a core model pre-trained with ample historical records, then fine-tuning for different tasks in the same city, analogous to a language model being fine-tuned for more specialized tasks.
@@ 6.4. Realistic Synthesis of Individual Events
To generate individual point data, we use our best performing model GRU to disaggregate from PUMA to BLOCK with NYC taxi data. Then, we measure the mutual information and KL divergence (using Python implementations444https://scikit-learn.org/stable/modules/generated/sklearn.metrics.normalized_mutual_info_score.html555https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kstest.html) between the distribution of disaggregated values and the original distribution in the same time period and region set. Then, for each BLOCK region with taxi counts, we randomly draw points within the region. We work with a set of block regions in Manhattan as shown in Figure 6. We draw samples from this set of regions for the on 12PM-1PM, 06/01/2016.
@@ 7. Discussion and Conclusions
We proposed a GRU model, which replaces simple matrix multiplications with spatial attention layers in the GRU model. The spatial attention layers model the complex spatial interactions among regions, while the gated recurrent module captures temporal dependencies. To further capture spatial interactions, we enforce containment relationships between layers found in practice by multiplying the spatial attention result by a containment mask ensuring that child region counts always sum to parent region counts. Our experimental results on both taxi and bikeshare dataset consistently demonstrated the effectiveness of our proposed GRU model, with statistically significant enhancements observed in almost all disaggreagtion tasks with faster convergence.
@@
Since not all variables are available with significant training data, we considered the transferability of a model trained on one variable for predictions on another (in the same city). Surprisingly, with only a week of fine-tuning bikeshare data, the taxi model exhibited performance comparable to, sometimes better than, that of the bikeshare-only model. This result exposes that the underlying dynamics of human activity may appear in all variables observed from the same built environment, motivating new applications that integrate multiple data sources to overcome limited data availabilityĀ (Khatiwada etĀ al., 2022). We also showed that a simple samplign approach can synthesize individual events from aggregate data, exposing new opportunities for research.
@@
There are some limitations to the study that we plan to address in future research: 1) we only considered the geographic levels for which containment holds. However, our spatial attention calculation is not limited by the containment property. We can still construct the containment maps even when a finer region overlap with more than one coarse regions. However, the extent to which these containment maps enhance learning effectiveness remains to be investigated in future studies. 2) Our current transfer learning experiments exclusively focus on scenarios where datasets originate from the same city and domain. This assumption implies shared geographic structures/grids and similar underlying dynamics. However, in practice, A) two variables could stem from different domains within the same city, resulting in varied underlying dynamics despite sharing the same city structure; B) transfer learning across cities, where the underlying dynamics may be different, is an important use case that we do not consider, and 3) We only investigated one set of regions to synthesize individual points for one time period. The spatial and temporal variances might impact the quality of the synthesized points. Additionally, we did not investigate whether the ability to synthesize individual events introduces privacy risk.

### v4

@@ 6.2. SARN Improves Neural Results
From Table 2, we observe that our proposed SARN has the best performance across all six tasks! With taxi data, it delivers a 5.0% improvement over the best neural models (GRU) on average and a 40.1% improvement over the best classical methods (HR) on average. With bikeshare data, we see relatively smaller but still notable improvements of 1.2% and 13.6%, respectively. The largest improvement is observed when disaggregating from PUMA to NTA with NYC Taxi data, where SARN reduces the average hourly taxi count error in the entire Manhattan region by 301 compared to the GRU model! For statistical significance tests, we use bootstrap sampling for 10,000 iterations (Jurafsky and Martin, 2008; Sonderegger and Buscaglia, 2020): On each iteration, we sample the predictions, then calculate error and compute the proportion that are better than the baseline to provide an empirical estimate of the p-value. We use 0.05 as the significance level. With both the taxi and bikeshare dataset, SARN consistently outperforms GRU with statistically significant improvement.
We notice that STT models generally perform poorly on both datasets, even worse than FNN models. The performance gap is noticeably large on simpler tasks, from PUMA to NTA. We conjecture two reasons: 1) STT has overly complicated architectures for simpler disaggregation tasks. STT has over 150K parameters when disaggregating PUMA to NTA, while GRU only has 10K and SRAN has 35K. Models with more parameters do not necessarily enhance performance, but could possibly contain noisy signals that hurt the results. 2) STT shares the same spatial transformer layer as SARN. The only difference is the temporal module. Temporal attention is not effective and is incapable of producing consistent results compared with recurrent modeling unit in the disaggregation task. Although attention-based solutions are increasingly popular due to their success, we find that relatively short sequences and smaller capacity models can still motivate recurrent models.
SARN Converges Faster than Other Neural Model:
@@
Figure 5 shows the convergence plots of neural models on two datasets. We notice that other than the disaggregation from PUMA to TRACT on NYC Bikeshare dataset, SARN models always start with, and maintain lower validation error throughout training stage, thus converging faster than other models. One explanation are — 1) given the same recurrent temporal module, SARN converges faster than GRU, indicating the power of spatial transformer layer. 2) The spatial attention layers in SARN models are aware of the spatial structures due to the containment maps, which provide ”hints” during training, steering the learning toward the correct solutions (which always exhibit the containment property), thereby converging faster.
SARN Yields Better Disaggregation Results (Spatial Investigation):
To facilitate understanding of the error distribution, we disaggregated NYC taxi data from PUMA to NTA using four neural models. Analyzing TRACT and BLOCK levels is challenging due to the large number of regions involved (283 and 3733 respectively). MAE error is calculated for all test samples and averaged for each NTA region. To compare performance, we calculate the difference in MAE error of each region between the other neural models and SARN, referred to as relative error. A positive relative error indicates that the neural model performs worse than SARN, while a negative relative error indicates better performance. Figure 6 visualizes these errors. We observe the following: 1). SARN tends to have higher errors in the central region of Manhattan. This is reasonable given the higher traffic volume in the city center, making accurate disaggregation more challenging. 2) Overall, the other three neural models produce larger errors compared to SARN (visualized as red regions). FNN and GRU generally have more errors in the central regions but perform similarly or slightly better in the upper regions. STT, however, makes more errors across all regions. In conclusion, SARN consistently disaggregates data more accurately than the other neural models.
Resolutions | Model | SARN | G.T. | % Diff |
NTA | LSTM | 184.665 | 184.035 | 0.03 |
AGCRN | 124.911 | 137.625 | -9.24 | |
STSGCN | 50.441 | 50.134 | 0.61 | |
TRACT | LSTM | 23.945 | 23.973 | -0.11 |
AGCRN | 9.764 | 9.808 | -0.45 | |
STSGCN | 10.586 | 10.145 | 4.35 | |
BLOCK | LSTM | 2.328 | 2.393 | -2.72 |
AGCRN | 1.648 | 1.452 | 13.47 | |
STSGCN | - | - | - | |
Average % Diff | 0.78 |
6.3. Disaggregated Data is Authentic and Benefits Downstream Application
Table 3 presents the next-hour traffic demand prediction results in MAE using disaggregated data from SARN at three geographic levels. These results are compared to those from models trained with ground truth data. We observe that, all models trained with disaggregated data exhibit performance that is very close to those trained with ground truth data, with minor and sometimes negligible differences. In some cases, the models perform even better with disaggregated data (e.g., NTA with AGCRN; TRACT with AGCRN; BLOCK with LSTM). Overall, the percentage difference is only 0.78% on average, which is very trivial. These findings indicate that the disaggregated values produced by the SARN model are authentic, of high quality, and closely mimic the distribution of the ground truth data. This has significant real-world implications. Disintegrating aggregated data into finer resolutions and then forecasting traffic demand can capture hotspots and dynamic human mobility patterns that coarse, aggregated data cannot provide.
@@ 6.4. Transfer Learning is Remarkably Effective
We trained SARN models on NYC taxi data and evaluated on bikeshare in the same NYC domain. We fine-tuned with one day (24 samples), 3 days (72), one week (168), two weeks (336), and one month of data (720). Figure 7 shows the results: adding more fine-tuning samples enhances disaggregation performance. Surprisingly, fine-tuning with only one month of data achieves results comparable with the specialized model trained from scratch with bikeshare data. Additionally, even the performance slightly dropped, the benefit gain in saving computational costs is significant — fine-tuning a pre-trained model greatly reduces the time costs (12 hours v.s. 12 hours). This result has significant implications for practical settings: Training data is not uniformly available for all variables due to technical, privacy, and political effects. But because cities are complex systems, all variables are interdependent (Howe et al., 2022). This result shows that pre-training can capture city dynamics resulting from the built environment and human activity, and that those signals tend to be predictive for multiple variables. We envision developing a core model pre-trained with ample historical records, then fine-tuning for different tasks in the same city, analogous to a language model being fine-tuned for more specialized tasks.
@@ 6.5. Realistic Synthesis of Individual Events
To generate individual point data, we use SARN to disaggregate from PUMA to BLOCK with NYC taxi data. Then, we measure the mutual information and KL divergence (using Python implementations555https://scikit-learn.org/stable/modules/generated/sklearn.metrics.normalized_mutual_info_score.html666https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kstest.html) between the distribution of disaggregated values and the original distribution in the same time period and region set. Then, for each BLOCK region with taxi counts, we randomly draw points within the region. We work with a set of block regions in Manhattan as shown in Figure 8. We draw samples from this set of regions for the on 12PM-1PM, 06/01/2016.
@@ 7. Discussion and Conclusions
We proposed a structurally-aware recurrent network SARN model, which replaces fully-connected layers with structurally-aware spatial attention (SASA) layers in the GRU model. The spatial attention layers model the complex spatial interactions among regions, while the gated recurrent module captures temporal dependencies. Each SASA layer calculates global attention that captures broad spatial dependencies. It also calculates structural attention using the containment maps, ensuring a coherent and consistent disaggregation across different geographic resolutions. Our experimental results on both taxi and bikeshare dataset consistently demonstrated the effectiveness of our proposed SARN model, with statistically significant enhancements observed in almost all disaggreagtion tasks. It also converges faster than other neural models.
@@
We conducted three supplementary experiments to demonstrate the utility of the disaggregated data — 1) We use the disaggregated data in traffic demand prediction task, and show that the disaggregated data closely approximately the ground truth data with trivial performance difference, indicating its great quality and potential applicability in downstrea applications where coarse, aggregated data is less useful. 2) Since not all variables are available with significant training data, we considered the transferability of a model trained on one variable for predictions on another (in the same city). Surprisingly, with only a month of fine-tuning bikeshare data, the model exhibited performance comparable to that of the specialized model. This result exposes that the underlying dynamics of human activity may appear in all variables observed from the same built environment, motivating new applications that integrate multiple data sources to overcome limited data availability (Khatiwada et al., 2022). 3) We show that a simple sampling approach can synthesize individual events from aggregate data, exposing new opportunities for research.
@@
There are some limitations to the study that we plan to address in future research: 1) we only considered the geographic levels for which containment holds. However, our spatial attention calculation is not limited by the containment property. We can still construct the containment maps even when a finer region overlap with more than one coarse regions. However, the extent to which these containment maps enhance learning effectiveness remains to be investigated in future studies. 2) Our current transfer learning experiments exclusively focus on scenarios where datasets originate from the same city and domain. This assumption implies shared geographic structures/grids and similar underlying dynamics. However, in practice, A) two variables could stem from different domains within the same city, resulting in varied underlying dynamics despite sharing the same city structure; B) transfer learning across cities, where the underlying dynamics may be different, is an important use case that we do not consider, and 3) We only investigated one set of regions to synthesize individual points for one time period. The spatial and temporal variances might impact the quality of the synthesized points. Additionally, we did not investigate whether the ability to synthesize individual events introduces privacy risk. 4) Given the sequential processing nature of GRU model, SARN does incur heavy computational costs compared to other baseline models. For example, when disaggregating from PUMA to BLOCK with NYC datasets, it could take 12 hours of training, while FNN or GRU only a few hours. The computational burden is especially heavy when the disaggregated dimension is high (e.g., block level). We consider this as a limitation of using sequential models.